a triangular sail on a sailboat is 6 meters across the bottom and 10 meters tall.what is the area of this sail
Answer by macston(5194)

For triangle Area-1/2(b)(h)
Area =1/2(6 m)(10 m)=1/2(60 sq m)=30 sq m
Area of the sail is 30 sq meters